# A spin on Hagedorn temperatures and string stars

### v1

@@ 6 Future directions
There are several natural generalizations of our work. One is to include more than one angular momentum (for $d>3$), and compare with the appropriate (Myers–Perry) black hole. Specifically for $d=4$, one can turn on equal angular momenta for the $1,2$ and the $3,4$ directions, in which the string star equations are one-dimensional again (as in $d=2$). It would also be interesting to better understand the AdS rotating solutions, especially for pure NS-NS, where one could hope for an exact worldsheet description following Agia:2023skp . In AdS, it is also common to turn on chemical potentials for rotations of internal direction Chamblin:1999tk (dual on the CFT to global symmetries). Looking for an AdS string star rotating in those directions will generalize the charged solutions of Chen:2021dsw .
In Harmark:2021qma , an integrability analysis of $\mathcal{N}=4$ SYM allowed for the computation of the CFT Hagedorn temperature even at strong coupling, with an exact matching with the bulk $l_{s}/l_{ads}$ expansion. Using the same methods, the authors studied the Hagedorn temperature in the presence of chemical potentials for angular momentum of the type (101). It would be illuminating to compare the strong coupling regime with (109). Following the recent progress Bigazzi:2023oqm ; Ekhammar:2023glu ; Bigazzi:2023hxt ; Ekhammar:2023cuj ; Harmark:2024ioq ; Bigazzi:2024biz ; Bigazzi:2024sjy ; Canneti:2025rsp ; Castellano:2025ljk one could study higher $l_{s}/l_{ads}$ orders of the Hagedorn temperature, either $R_{H}(\Omega)$ (109) or $R_{H}(\nu)$ (108). Following the analysis of section 5, it is also possible to study the Hagedorn temperature and rotating winding strings condensates in the presence of an (holographic) angular twist $\nu$ in confining backgrounds Urbach:2023npi .
More conceptually, we hope for a better understanding of the rotating string star solution at large angular momentum, and its possible transition to the quantum string rod Ceplak:2023afb . One could also wonder if there is a string star version for the extremal black hole, in the double limit of zero temperature and $J\sim g_{s}^{-\frac{4}{d-1}}$ (at which $r_{h}\sim l_{s}$).
@@ A.1 Hagedorn temperatures
We now turn to the spin $1$ fields, beginning with the complex boson $\Phi=X^{1}+iX^{2}$. Diagonalizing $H,P,$ and $J$ together gives the eigenvalues $J=\pm 1$ for each oscillator Russo:1994ev ; Matsuo:2009sx . The complex boson partition function of a single chirality gives
@@
where in the second line we use the modular properties of $\vartheta_{11},\eta$ Polchinski:1998rq . To find the high-temperature behavior, we set $b=-2\pi i\tau$ and focus on the exponential form in the $b\ll 1$ limit. For brevity, we will assume the fundamental region $1/2<\nu<1/2$. The $\eta$ and $\vartheta_{11}$ functions satisfy in the $e^{2\pi i\cdot-1/\tau}=e^{-4\pi^{2}/b}\rightarrow 0$ limit the identities
@@
Moving to type II string theory, the new component is, of course, from the GSO-projected worldsheet fermions. The GSO projection is best understood in the case of periodic boundary conditions for the target space fermions. While $\nu=0$ stands for the anti-periodic boundary conditions, the periodic case is exactly $\nu=1$. We therefore write down the standard GSO-projected partition function Polchinski:1998rr , only with angular velocity $\hat{\nu}=1-\nu$ ($1<\nu<1$) (see Ferko:2024uxi for a recent discussion)
@@ A.2 Entropy at fixed $J/M^{2}$
In this section, we would like to study the microcanonical entropy $S(M,J)$ in the scaling $M^{2}\gg 1$ for fixed $J/M^{2}$, mostly following Matsuo:2009sx . To do that, we begin by writing down the worldsheet partition function in the high temperature limit $b\ll 1$ for a given chirality. For any of the string theories we described, it has the form
@@ A.3 The $J/M$ expansion
Here, we derive some formulas for the free string in the scaling $M^{2}\gg 1$, with fixed $J/M\ll 1$, following Russo:1994ev . The starting point is the same as before: the worldsheet canonical partition function $Z(b,\nu)$. We focus first on the left chirality, and then combine it with the right chirality. As we will see, in terms of the worldsheet parameters, we take the limit of constant $\nu/b$ and small $b$. In the language of Russo:1994ev , this corresponds to the case of constant $\lambda:=2\pi\nu/b$. The asymptotic level density for bosonic string theory in that case is Russo:1994ev
@@
where $\beta_{H,L}=4\pi$ is the contribution of the left chirality to the bosonic ($\nu=0$) Hagedorn temperature. The power of the prefactor has two contributions: the 24=26224=26-2, which is the bosonic central charge in light-cone quantization, and the (theory-independent) $5$, which comes from the Cardy formula and the $J$-measure Russo:1994ev . Expanding (142) in $j_{L}/\sqrt{N_{L}}$ gives the asymptotic density of states
@@
It is straightforward to write down the asymptotic density for a $\mathcal{N}=2$ SUSY chirality Russo:1994ev
@@
Following Ceplak:2024dxm , we estimate the sizes by taking averages of the worldsheet position for different axes. For the size orthogonal to the plane of rotation $L_{\perp}^{2}$, we compute

### v2

@@ 6 Future directions
There are several natural generalizations of our work. One is to include more than one angular momentum (for $d>3$), and compare with the appropriate (Myers–Perry) black hole. Specifically for $d=4$, one can turn on equal angular momenta for the $1,2$ and the $3,4$ directions, in which the string star equations are one-dimensional again (as in $d=2$). It would also be interesting to better understand the AdS rotating solutions, especially for pure NS-NS, where one could hope for an exact worldsheet description following Agia and Jafferis (2023). In AdS, it is also common to turn on chemical potentials for rotations of internal direction Chamblin et al. (1999) (dual on the CFT to global symmetries). Looking for an AdS string star rotating in those directions will generalize the charged solutions of Chen et al. (2023).
In Harmark and Wilhelm (2022), an integrability analysis of $\mathcal{N}=4$ SYM allowed for the computation of the CFT Hagedorn temperature even at strong coupling, with an exact matching with the bulk $l_{s}/l_{ads}$ expansion. Using the same methods, the authors studied the Hagedorn temperature in the presence of chemical potentials for angular momentum of the type (101). It would be illuminating to compare the strong coupling regime with (109). Following the recent progress Bigazzi et al. (2023b); Ekhammar et al. (2023b); Bigazzi et al. (2023a); Ekhammar et al. (2023a); Harmark (2024); Bigazzi et al. (2024, 2025); Canneti et al. (2025); Castellano and Zatti (2025) one could study higher $l_{s}/l_{ads}$ orders of the Hagedorn temperature, either $R_{H}(\Omega)$ (109) or $R_{H}(\nu)$ (108). Following the analysis of section 5, it is also possible to study the Hagedorn temperature and rotating winding strings condensates in the presence of an (holographic) angular twist $\nu$ in confining backgrounds Urbach (2023).
More conceptually, we hope for a better understanding of the rotating string star solution at large angular momentum, and its possible transition to the quantum string rod Čeplak et al. (2023). One could also wonder if there is a string star version for the extremal black hole, in the double limit of zero temperature and $J\sim g_{s}^{-\frac{4}{d-1}}$ (at which $r_{h}\sim l_{s}$).
@@ A.1 Hagedorn temperatures
We now turn to the spin $1$ fields, beginning with the complex boson $\Phi=X^{1}+iX^{2}$. Diagonalizing $H,P,$ and $J$ together gives the eigenvalues $J=\pm 1$ for each oscillator Russo and Susskind (1995); Matsuo (2010). The complex boson partition function of a single chirality gives
@@
where in the second line we use the modular properties of $\vartheta_{11},\eta$ Polchinski (2007a). To find the high-temperature behavior, we set $b=-2\pi i\tau$ and focus on the exponential form in the $b\ll 1$ limit. For brevity, we will assume the fundamental region $1/2<\nu<1/2$. The $\eta$ and $\vartheta_{11}$ functions satisfy in the $e^{2\pi i\cdot-1/\tau}=e^{-4\pi^{2}/b}\rightarrow 0$ limit the identities
@@
Moving to type II string theory, the new component is, of course, from the GSO-projected worldsheet fermions. The GSO projection is best understood in the case of periodic boundary conditions for the target space fermions. While $\nu=0$ stands for the anti-periodic boundary conditions, the periodic case is exactly $\nu=1$. We therefore write down the standard GSO-projected partition function Polchinski (2007b), only with angular velocity $\hat{\nu}=1-\nu$ ($1<\nu<1$) (see Ferko et al. (2025) for a recent discussion)
@@ A.2 Entropy at fixed $J/M^{2}$
In this section, we would like to study the microcanonical entropy $S(M,J)$ in the scaling $M^{2}\gg 1$ for fixed $J/M^{2}$, mostly following Matsuo (2010). To do that, we begin by writing down the worldsheet partition function in the high temperature limit $b\ll 1$ for a given chirality. For any of the string theories we described, it has the form
@@ A.3 The $J/M$ expansion
Here, we derive some formulas for the free string in the scaling $M^{2}\gg 1$, with fixed $J/M\ll 1$, following Russo and Susskind (1995). The starting point is the same as before: the worldsheet canonical partition function $Z(b,\nu)$. We focus first on the left chirality, and then combine it with the right chirality. As we will see, in terms of the worldsheet parameters, we take the limit of constant $\nu/b$ and small $b$. In the language of Russo and Susskind (1995), this corresponds to the case of constant $\lambda:=2\pi\nu/b$. The asymptotic level density for bosonic string theory in that case is Russo and Susskind (1995)
@@
where $\beta_{H,L}=4\pi$ is the contribution of the left chirality to the bosonic ($\nu=0$) Hagedorn temperature. The power of the prefactor has two contributions: the 24=26224=26-2, which is the bosonic central charge in light-cone quantization, and the (theory-independent) $5$, which comes from the Cardy formula and the $J$-measure Russo and Susskind (1995). Expanding (142) in $j_{L}/\sqrt{N_{L}}$ gives the asymptotic density of states
@@
It is straightforward to write down the asymptotic density for a $\mathcal{N}=2$ SUSY chirality Russo and Susskind (1995)
@@
Following Čeplak et al. (2024), we estimate the sizes by taking averages of the worldsheet position for different axes. For the size orthogonal to the plane of rotation $L_{\perp}^{2}$, we compute